OpenStudy (anonymous):

3/4x+1 find all numbers for which the rational expression is undefined

OpenStudy (anonymous):

x/3=4/x+1 x/3 = 4/ (x+1) x(x+1)=12 x^2+x-12=0 x^2+4x-3x-12=0 x(x+4)-3(x+4)=0 (x-3)(x+4)=0 x=3 OR -4

OpenStudy (anonymous):

can i get a woot woot ! - <3 seashell

OpenStudy (anonymous):

lol

OpenStudy (asnaseer):

is it?:\[\frac{3}{4x+1}\]

OpenStudy (asnaseer):

if so, this is undefined when the denominator is zero, i.e. when:\[4x+1=0\]